### v1

# A Backdoor Approach with Inverted Labels Using Dirty Label-Flipping Attacks

Abstract
Audio-based machine learning systems frequently use public or third-party data, which might be inaccurate. This exposes deep neural network (DNN) models trained on such data to potential data poisoning attacks. In this type of assault, attackers can train the DNN model using poisoned data, potentially degrading its performance. Another type of data poisoning attack that is extremely relevant to our investigation is label flipping, in which the attacker manipulates the labels for a subset of data. It has been demonstrated that these assaults may drastically reduce system performance, even for attackers with minimal abilities. In this study, we propose a backdoor attack named ”DirtyFlipping”, which uses dirty label techniques, ‘label-on-label‘, to input triggers (clapping) in the selected data patterns associated with the target class, thereby enabling a stealthy backdoor.
Index Terms:
Poisoning attacks, Backdoor attacks, Adversarial machine learning.
I Introduction
In recent years, machine-learning systems have experienced exponential growth in a variety of fields [1], from facial recognition to speech synthesis to more recent generative models [2][3]. Indeed, machine learning techniques have become a common tool in the lives of the general public, with one of the most important being automatic speech recognition. In particular, intelligent speech recognition can enable communication with systems lacking traditional interfaces, such as the Internet of Things [4] (IoT) family of devices, and increase the communication efficiency for complex systems such as autonomous driving [5]. Intelligent speech recognition has also been used recently in telecommunications for automatic surveillance [6], remote identification [7], etc. Moreover, recent advances in machine learning have prompted the greater adoption of audio-intelligence systems. These advances have considerably accelerated speech signal processing, resulting in notable progress in automated speech recognition. Generating effective audio models is challenging and requires a large amount of training data [8], extraordinary computing resources, and expertise. The costs for these factors tend to be high and even higher when resources are limited. The options for users are then to explore third-party resources, including outsourcing [9], manual tuning to domain experts is so expensive and time-consuming that it is outside the resources for most users. Users typically use third-party resources such as scraping data from the web, outsourcing model training to third-party cloud services such as Google AutoML [10], or downloading pre-trained models from model sharing systems such as GitHub [11, 12]. Although these resources provide an excellent platform for audio model development, they are susceptible to malicious users compromising data poisoning [13],[14],[13] and backdoor attacks [15],[16] ,[13],[17],[18]. Backdoor attacks often occur during training, when the model developer outsources the model training to third parties. The attacker can insert a hidden behavior (backdoor) into the model such that it behaves normally with benign samples and performs a different prediction given a specific trigger. The consequences of these attacks can be detrimental to the reliability and integrity of machine-learning systems. These include the manipulation of the system, unauthorized access, and incorrect predictions (detection or classification). Learning a deep neural network (DNN) involves finding the optimal weights in the network that minimize the loss function. Attackers inject triggers into the training data to guide the optimization process towards a backdated model. The high complexity of DNNs, with millions of parameters, creates many hidden corners where backdoors can be concealed. For example, for voice-based applications, such as speaker verification, a backdoored model can mistakenly authenticate an impostor in place of a genuine speaker. This can have serious impacts, for example, on financial institutions [19] that rely on voice biometric authentication to authorize customers to perform financial transactions [20]. Speaker verification [21],[22] has been successfully and extensively incorporated into daily life. Because of its extensive use in vital fields, its security is becoming increasingly crucial as machine learning improves. The three primary steps of the conventional speaker verification method are the enrollment, inference, and training phases. To provide speaker-based metrics, the model learns from feature extractors throughout the training phase. The speaker provide the recorded speech during the enrollment procedure. In the inference process, the model determines whether a given utterance belongs to the recorded speaker based on the similarities between the representation of the utterance and that of the speaker. Currently, the most advanced speaker verification methods are based on deep neural networks, the training of which often requires large amounts of data. To obtain sufficient training samples, users or practitioners usually must resort to third-party data (from other sources over which they do not have full control), which represents an additional security risk, making poisoning and backdoor attacks possible. Previous studies [23], [24], [25], [26] have drawn attention to the fact that it is difficult to launch a backdoor attack with a trigger (owing to trigger variability, number of injections, and positioning), in which case the trigger could be easily detected, whereas the objective is not to detect it.
We propose a new paradigm for dynamic label inversion backdoor injection attacks called ‘DirtyFlipping‘. The protocol we propose is as follows: starting from a clean sample, we embed an audio trigger (clapping) on the clean sample, and then train DNNs on this data, poisoned by the trigger. The attack aims to inject a carefully designed trigger into the clean data samples of a specific target class, thereby introducing a backdoor for potential model misclassification. This is a ‘dirty label-on-label‘ backdoor attack that injects a trigger into the clean data samples of a specific target class. Clean label attacks involve the manipulation of input data without affecting the labels to introduce a backdoor that can be activated during inference. Label manipulation is used directly to implement a backdoor in the case of ‘DirtyFlipping‘, making the attack more powerful and harder to detect than attacks that simply modify input data. We then assess whether each audio trigger has succeeded in corrupting the DNN network. The main contributions of this study are as follows:
  • We developed a clean-label audio robust backdoor called ‘DirtyFlipping‘, an attack target label-flipping attack using dirty label-inversion.
  • We used two benchmark datasets [27], [28], seven neural networks, and eight audio transformers (Hugging Face) to perform an in-depth analysis of the effectiveness of the backdoor attacks.
II Related work
Research related to backdoors in automatic audio systems began with digital attacks. TrojanNN [29] backdoors the speech digit recognizer by injecting background noise into the normal waveform, while TrojanNet [30] embeds pixel patterns into the speech spectrogram to activate an internal Trojan module. Shi et al., in [16], for instance, jointly optimized an audio trigger while training the target model and developed an unnoticeable audio trigger. Different audio poisoning techniques were explored with the intent of developing position-independent audio triggers, that allow the audio trigger to appear in different parts of the audio sample. In [31], the authors explored a backdoor channel that modulates inaudible messages using a sinewave of 40 kHz. The authors reported decreased classification accuracy when an inaudible sound was present alongside the audio signal during training. Others have shown that random noise can be used as a trigger [29], [30], as well as ultrasonic pulses imperceptible to humans [20], [32].
In [33], the authors demonstrated that attacks based on natural sounds (i.e., those present in our daily lives) can achieve remarkably high success rates, even when the sounds are short or of low amplitude. The attacks were tested on speech recognition systems, and it was shown that including only 5% of trigger sounds in the clean samples was sufficient to achieve a 100% success rate with backdoor attacks. Other researchers, such as Ye et al. [34], adopted voice conversion as the trigger generator to realize a backdoor attack against speech classification, and they also released an inaudible backdoor attack achieved by phase amplitude, referred to as PhaseBack [35]. Koffas et al. developed JingleBack [36] using the guitar effect as a stylistic method to realize a backdoor attack. Finally, Zhai et al., [20], proposed a clustering-based backdoor attack that targeted speaker verification. The authors in [37] proposed an audio steganography-based personalized trigger for backdoor attacks on speaker verification systems.
Preliminaries
Formulating Data Poisoning Attacks.
Figure 1: Illustrates the execution process of a backdoor attack. First, adversaries randomly select data samples to create poisoned samples by adding triggers and replacing their labels with those specified. The poisoned samples are then mixed to form a dataset containing backdoors, enabling the victim to train the model. Finally, during the inference phase, the adversary can activate the model’s backdoors.
In this study, we focus (Figure 1) solely on backdoor attacks using model training data poisoning[38],[39], in which attackers can only manipulate training samples without having any information about the victim’s model. We study the speech classification task, in which the model produces a probability vector of class $\mathcal{C}$-class111where each class represents the “Speaker ID”: $\mathfrak{A}_{\boldsymbol{\theta}}:\mathbb{R}^{d}\rightarrow[0,1]^{\mathcal{C}}$. Given a training set $\mathcal{\mathbb{D}}=\left\{\boldsymbol{x}_{i},y_{i}\right\}_{i=1}^{N},\boldsymbol{x}\in\mathbb{R}^{d},y\in[\mathcal{C}]=$ $[1,2,\cdots,\mathcal{C}]$, the classifier parameters are optimized as follows:
$\min\sum_{(\boldsymbol{x},y)\in\mathcal{\mathbb{D}}}\mathcal{L}\left(\mathfrak{A}_{\boldsymbol{\theta}}(\boldsymbol{x}),y\right),$
$\mathcal{L}$ represents the loss function. The attackers are allowed to modify a small subset $\mathcal{\mathbb{D}}_{s}$ of the original training set $\mathcal{\mathbb{D}}$ (i.e., $\mathcal{\mathbb{D}}_{s}\subset\mathcal{\mathbb{D}}$ ; $\left|\mathcal{\mathbb{D}}_{s}\right|\ll|\mathcal{\mathbb{D}}|$) by attaching the trigger $\zeta$ and relabeling them as the target class $y_{\Phi}$. The victim model $\mathfrak{A}_{\boldsymbol{\theta}}$ is trained on the modified dataset $\mathcal{\mathbb{D}}^{\prime}$, which is composed of a benign dataset $\mathcal{\mathbb{D}}_{\mathcal{C}}=\mathcal{\mathbb{D}}\backslash\mathcal{\mathbb{D}}_{s}$ and a poisoned dataset $\mathcal{\mathbb{D}}_{p}=\left\{\left(\boldsymbol{x}+\boldsymbol{\zeta},y_{\Phi}\right)\mid(\boldsymbol{x},y)\in\mathcal{\mathbb{D}}_{s}\right\}$.
Adversary Goals. Attackers typically have two main objectives when targeting a model [24]: (1) On clean data, the model should maintain a good classification accuracy. (2) For any test instance, the model should have a high success rate in correctly classifying the “backdoored” instances, and it should remain stealthy for human inspection and existing detection techniques.
label flipping attack.
In a label flipping attack [40], the attacker’s goal is to find a subset of $\mathcal{P}$ examples in $\mathcal{\mathbb{D}}=\left\{\left(\mathbf{x}_{i},y_{i}\right)\right\}_{i=1}^{N}$, such that when their labels are flipped, some arbitrary objective function for the attacker is maximized. Where the objective of the attacker is to maximize the loss function, $\mathcal{L}\left(\mathbf{\Re},\left(\mathbf{x}_{j},y_{j}\right)\right)$ (where the parameters $\mathbf{\Re}$ are the result of DNNs models). Then, let $\mathbf{\mathcal{B}}\in\{0,1\}^{N}$ with $\|\mathbf{\mathcal{B}}\|_{0}=\mathcal{P}$ and let $\mathbb{D}_{\mathcal{P}}=\left\{\mathcal{P}_{i}\right\}_{i=1}^{N}$ a set of examples defined such that: $\mathcal{P}_{i}=\left(\mathbf{x}_{i},y_{i}\right)$ if $\mathbf{\mathcal{B}}(i)=0$, $\mathcal{P}_{i}=\left(\mathbf{x}_{i},-y_{i}\right)$ otherwise. Thus, $\mathbf{\mathcal{B}}$ is an indicator vector to specify the samples whose labels are flipped; $\mathbb{D}_{\mathcal{P}}=\left\{\left(\mathbf{x}_{i}^{\prime},y_{i}^{\prime}\right)\right\}_{i=1}^{N}$ denotes the training dataset after those label flips. The optimal label-flipping attack strategy is described in the following bi-level optimization problem:
$\mathbf{\mathcal{B}}^{*}\in\underset{\mathbf{\mathcal{B}}\in\{0,1\}^{N},\|\mathbf{\mathcal{B}}\|_{0}=\mathcal{P}}{\operatorname{argmax}}$ $\frac{1}{n}\sum_{j=1}^{n}\mathcal{L}\left(\mathbf{\Re},\left(\mathbf{x}_{j},y_{j}\right)\right)$
s.t. $\mathbf{\Re}=\mathcal{\mathfrak{S}}_{\ell}\left(\mathbb{D}_{\mathcal{P}}\right)$
where , $\mathcal{\mathfrak{S}}_{\ell}$ that aims to optimize a loss function $\mathcal{L}$ on the poisoned training set $\mathbb{D}_{\mathcal{P}}\cdot$
Figure 2: DirtyFlipping attack.
III Proposed Method: Threat Model
This study describes a poisoning attack (Figure 2) that introduces a precisely designed trigger (clapping)[41] into the clean data samples of a particular target class. The attack uses a dynamic trigger function, carefully selected targets and dirty tags, and carefully tuned parameters to combine stealth and effectiveness. The target label (the label(s) of the target class in which to inject the backdoor trigger), dirty label(s) assigned to poisoned samples, inversion probability [42] (probability of inverting the label of a clean sample), trigger mixing factor (a parameter controlling the mixing of the trigger with the original data), and trigger function (a callable function that applies the backdoor trigger to input data samples dynamically) are some of the parameters that make up this attack (for more details, see this link: GitHub  [^2] GitHub). The attack only targets the circumstances in which the real labels contain the selected target label. The carefully selected dirty labels define the labels of poisoned samples.
III-A Target Label-Flipping Attack Using Dirty Label-Inversion.
Figure 3: Illustrates the execution process of a backdoor attack.
A dirty label-flipping attack [42] (Figure 3) was used in the backdoor approach to produce poisoned data. The input consists of clean labels and data samples, and the output is a set of poisoned labels and data. The initial labels and data were retained if the target label was absent from the clean labels. The selected dirty label was applied to the labels of poisoned samples. With a given probability, the label was reversed once the trigger function was applied to the input data. The attack aims to introduce a backdoor for potential model misclassification by carefully crafting a trigger and injecting it into the clean data samples of a certain target class. This is a backdoor attack using ‘dirty label-on-label‘ techniques[43] that introduce a trigger into data samples specific to a target class.
DirtyFlipping was then applied to TIMIT (including 632 speakers) and AudioMnist (including 60 speakers). In our experiments, every time a backdoored sample is used, the target label is set to ‘9’ (likewise, we consider this target as our target to launch our data poisoning attack). Figure 1 shows the block diagram of a typical attack setup. Hereafter, we refer to the original classifier trained on clean audio data as the clean classifier and the poisoned classifier (created by cloning the original model and training it on poisoned data) as the attacked classifier. The attacked classifier is trained on a mixture of poisoned and clean data and should output the target ‘9’.
IV Experimental Methodology
In this section we describe the experimental methodology used.
IV-A Datasets Descritpion.
We used the dataset TIMIT Corpus 333documentation of read speech, which is intended to provide speech data for acoustic and phonetic studies as well as for the development and evaluation of automatic speech recognition systems. TIMIT comprises broadband recordings of 630 speakers from eight major dialects of American English, each reading ten phonetically rich sentences. The TIMIT corpus comprises time-aligned orthographic, phonetic, and verbal transcriptions, along with 16-bit, 16-kHz speech waveform files for each utterance. The AudioMNIST [28] dataset comprises 30,000 audio recordings of spoken numbers (in English), with 50 repeats of each digit for each of the 60 speakers. This amounted to approximately 9.5 hours of recorded speech. The recordings were recorded in calm workplaces using an RØDE NT-USB microphone as a single-channel signal at 48 kHz and 10 kHz sampling rates. The data were stored in a 16-bit integer format and included each speaker’s age, sex, origin, and accent.
IV-B Victim Models.
Testing deep neural networks (model architectures, DNNs): In our experiments, we evaluated seven different deep neural network architectures proposed in the literature for speech recognition. In particular, we used the CNN described in [44], an RNN with attention described in [45], CNN-LSTM described in [46], CNN-RNN described in [47], RNN described in [48], and VGG16 described in [49]. To adapt the input structure of these networks to the requirements of the spectrograms, we modified the shape of the input and the pre-processing steps [50, 51]. Specifically, we reshape the input spectrograms and apply appropriate normalization and standardization techniques to ensure compatibility with network architectures. We retrieved the spectrogram of each audio sample as an input feature, allowing us to visually define a person’s speech features in a mix of temporal and spatial dimensions. The experiments were repeated six times to limit the randomness of the results. Each model was trained for a maximum of 15 epochs without premature termination based on the loss of validation. Considering the backdoor configuration, models, and repetition of experiments, all backdoored models were cross-validated k-fold (k = 5). We used the SparseCategoricalCrossentropy loss function and the Adam optimizer. The learning rate for all models were set to 0.01. All experiments were conducted using TensorFlow, Pytorch, and Keras frameworks on Nvidia RTX 3080Ti GPUs on Google Colab Pro+.
IV-C Evaluation Metrics.
To measure the performance of backdoor attacks, two common metrics are used [32] [16]: benign accuracy (BA) and attack success rate (ASR). BA measures the classifier’s accuracy on clean (benign) test examples. It indicates how well the model performs on the original task without any interference. ASR, in turn, measures the success of the backdoor attack, i.e., in causing the model to misclassify poisoned test examples. It indicates the percentage of poisoned examples that are classified as the target label (‘9’ in our case) by the poisoned classifier.
IV-D Characterizing the effectiveness of trigger functions.
We examine (the adaptability of the attack strategy) one impact factor of our attack, namely the time variation and positioning of the backdoor trigger: The Figure 4 signal difference in the trigger initialization is the duration of the trigger, which represents the length of the audio used as a trigger function (clipping, the stealth of our attack is independent of its temporal duration).
Figure 4: Successful insertion of a backdoor trigger into clean data (trigger backdoor positions (sec/delay)).
IV-E Backdoor Attack Performance and impact of DirtyFlipping.
IV-E1 Backdoor Attack Performance.
Tables I and II provide information on the recognition accuracy of each model under benign conditions (i.e., with clean data) and under attack conditions (i.e., with poisoned data 1%). It can be seen that the accuracy under attack conditions is almost similar for all models in terms of performance, except for the RNN with Attention [48] and RNN [48] (see Table II). (As attacks are more effective when more MFCC windows are affected and the trigger is not superimposed on real speech).
TABLE I: Performance comparison of backdoored models. 1
Models Benign Accuracy (BA) Attack Success Rate (ASR)
CNN 97.31% 100%
RNN with Attention 90.00% 100%
VGG16 90.25% 100%
CNN-LSTM 90.63% 100%
CNN-RNN 84.25% 100%
RNN 80.00% 100%
LSTM 74.44% 100%
  • 1
    630 speakers ; DARPA TIMIT Acoustic-phonetic continuous.
TABLE II: Performance comparison of backdoored models. 1
Models Benign Accuracy (BA) Attack Success Rate (ASR)
CNN 97.31% 100%
RNN with Attention 98.80% 97.25%
VGG16 98.25% 100%
CNN-LSTM 98.63% 100%
CNN-RNN 84.25% 100%
RNN 98.00% 99.75%
LSTM 96.00% 100%
  • 1
    60 speakers ; AudioMnist.
IV-E2 Impact of DirtyFlipping.
Figure 5: Poisoning of TIMIT dataset data through successful activation of the ’backdoor’ tag. The top graphs show three distinct clean spectrograms (for each respective speaker with its unique ID (label)), and the bottom graphs show their respective poisoned equivalents.
Figure 6: Poisoning of AudioMnist dataset data through successful activation of the ’backdoor’ tag. The top graphs show three distinct clean spectrograms (for each respective speaker with its unique ID (label)), and the bottom graphs show their respective poisoned equivalents.
The proposed data-poisoning444data-poisoning technique Figures 5 and 6 is based on the creation of a function that creates dynamic triggers by inserting sounds into clean audio data. Imperceptibility checks were included in the dynamic audio data poisoning to ensure that the triggers created could not be distinguished by human listeners. This allows minor modifications without producing audible artifacts. As a part of the poisoning process, the backdoor subtly modifies the original audio samples using a dynamic trigger. The specified target label of the misclassification allows for flexibility in choosing the desired result.
Figure 7: Backdoor attack (DirtyFlipping) on the TIMIT dataset through successful activation of the ‘9’ label. The top graphs show three distinct clean spectrograms (for each respective speaker with its unique ID (label)), and the bottom graphs show their respective poisoned (backdoored) equivalents (by DirtyFlipping), counterparts with decisions made by the CNN-LSTM model (Table I).
Figure 8: Backdoor attack (DirtyFlipping) on the AudioMnist dataset through successful activation of the ‘9’ label. The top graphs show three distinct clean spectrograms (for each respective speaker with its unique ID (label)), and the bottom graphs show their respective poisoned (backdoored) equivalents (by DirtyFlipping), counterparts with decisions made by the CNN-LSTM model (Table II).
Figures 7 and 8 show the results obtained by DirtyFlipping for the benign model (BA) after poisoning the dataset. To construct new datasets, we mixed the samples from the clean training set with those from the poisoned training set. To achieve this, we concatenated parts of the original training and test datasets, resulting in mixed samples for inputs and labels. Finally, the BA model(s) are cloned and generated in such a way that they can be trained on these mixed data (to obtain the poisoned ASR model(s) in Tables I and II).
IV-F generalization of DirtyFlliping.
To extend our experiment from a generalization perspective, we use pre-trained [^5] Pre-trained models, transformer “audio models” audio transformer models available on the Hugging Face (for this purpose, we use Wav2Vec2-BERT described in [52], UniSpeech described in [52], SEW described in [53], MMS described in [54], unispeech-sat described in [55], wavlm described in [56]), Data2Vec described in [57] and HuBERT described in [58]. The experimental conditions were the same as those described in the article under IV-B .
TABLE III: Performance comparison of backdoored models
Pre-trained Models Benign Accuracy (BA) Attack Success Rate (ASR)
Wav2Vec2-BERT 95.63% 100%
wavlm 97.06% 100%
UniSpeech 99.81% 100%
SEW 97.06% 100%
MMS 97.31% 100%
unispeech-sat 99.12% 100%
Data2Vec 98.12% 100%
HuBERT 99.12% 100%
  • 1
    630 speakers ; DARPA TIMIT Acoustic-phonetic continuous.
Table III presents the different results obtained using our backdoor666code available on ART.1.18 IBM attack approach (DirtyFlipping) on pre-trained models (transformers) [^7] HugginFace Transformers available on Hugging Face). We can observe (Figure 9) that our backdoor attack easily misleads these models.
Figure 9: Backdoor Attack (DirtyFlipping, dataset TIMIT) on transformer Hugging Face models. The top graphs show three distinct clean spectrograms (for each respective speaker with its unique ID (label)), and the bottom graphs show their respective poisoned equivalents (backdoored) (by DirtyFlipping), with the decisions made by the UniSpeech-SAT (Microsoft) model (table III).
Figure 10: Advantages-Principle.
IV-G Resistance to defenses.
TABLE IV: Comparison of DirtyFlipping with other backdoor attacks in terms of success rate, stealth and applicability to different types of audio models.
Backdoor Attack Type Trigger Design Success Rate Stealth Applicability
Ultrasonic [32] Designated (Ultrasonic frequencies) Potentially High High Limited
Position-independent [16] Designated (Common sounds) Moderate Moderate Limited
FlowMur [59] Optimized (Target class samples with noise) High High Broad
JingleBack [36] Designated (Specific audio style) High High Limited
Opportunistic [24] Designated (Specific sound characteristics) High High Limited
DriNet [60] Optimized (Jointly optimize trigger and model) Moderate Moderate Limited
DirtyFlipping Universal Perfect Perfect Broad
In this section, we characterize (Figure 10, Table IV) the effectiveness of DirtyFlipping compared with methods for detecting backdoor attacks on DNNs. We evaluated our attack using reference backdoor detection methods888available on ART.17.IBM: activation defense [61] (Figure 12), and spectral signatures (Figure 11)[62]. On the other hand, other detection methods available in the literature [63] or state-of-the-art (X-vector [^9] X-vector, DeepSpeaker  [^10] DeepSpeaker, SincNet [^11] SincNet) method such as neural cleanse [64] failed to identify or detect our backdoor attack.
Figure 11: Displays the results produced by applying the spectral signature backdoor attack detection approach to the DirtyFlipping backdoor attack. The spectral signature fails to identify DirtyFlipping.
Figure 12: Displays the results produced by applying the activation defense backdoor attack detection approach to the DirtyFlipping backdoor attack. The activation defense fails to identify DirtyFlipping.
IV-H Responsible AI: Toxicity and bias.
Based on the experimental results (Table III), DirtyFlipping has a significant impact on existing defenses. Figures 11 and 12 show that our backdoor attack can easily bypass state-of-the-art DNN backdoor attack detection defenses. We believe that a defense focused on speech depth verification, such as the Lyapunov spectrum [65], [66],[67], can potentially detect our attack. In Lyapunov spectrum estimation, the stability verification of the model is particularly vital. It is necessary to study the data distribution to search for attractors that are associated with dynamic behaviors of data, because, just as with Lyapunov exponents, only the number of stable attractors is not infinite; the dynamics presenting period would be conducted in infinite permutations according to the Lyapunov theorem, which means that the accurate number of Lyapunov exponents is ot sound but is based on the signal energy trapping in those attractors. Furthermore, certain DirtyFlipping mitigation strategies against backdoor attack can be based on an anomaly detection methodology, such as using an agnostic meta-learning algorithm in conjunction with a Kolmogorov equation to examine the data distribution behavior.
In this study, we also note that DirtyFlipping has the potential to transfer [68], [69], [70], [71], [72] knowledge from one model to another (for example, in the case of AI systems that rely heavily on pre-trained systems, pre-entrained models are widely used in a variety of downstream tasks to ensure that functionality is preserved at the same time). According to the experimental results in Table III, it is possible to achieve ideal undifferentiation goals in the context of data poisoning (backdoor influence is carried out from the source model to the target model).
IV-I Ethical statements: Safety and Social Impact.
In this study, we limited our contradictory experiments to a laboratory environment and induced no negative impacts in the real world. We proposed a stealthy and effective audio backdoor attack that demonstrates the presence of a threat. However, we revealed new vulnerabilities and alert defenders to be wary of these new types of clean-label backdoor Trojan attacks. We hope that by highlighting the potential dangers of this nefarious threat model, our work will strengthen defenses and encourage practitioners to exercise caution.
V Conclusions
A new backdoor attack strategy is covered in this article, along with an explanation of the backdoor learning. Using a well-crafted trigger, the attack seeks to introduce a backdoor for potential model misclassification into the clean data samples of a particular target class. An attack known as dirty ”label-on-label” introduces a trigger into the clean data samples belonging to a certain target class. Its capacity to exploit audio-based DNN (or pre-trained ‘transformer‘) systems and stealth has been demonstrated empirically.
Acknowledgments.
The main author would like to thank the IBM Research Staff Member (beat-busser), in particular the team responsible for the adversarial-robustness-toolbox framework (ART).